OpenStudy (anonymous):

multiplication with roots: When you multiply a number by a simplified root: example: 5*6sqrt2 is that then 30sqrts2??

OpenStudy (anonymous):

if you mean \[\sqrt[2]{5*6}=\sqrt[2]{30}\] then you r right :)

OpenStudy (anonymous):

I mean: \[5\times (6\sqrt{2})\]

OpenStudy (anonymous):

\(30\sqrt2\)

OpenStudy (anonymous):

so I was right! :) Awesome. :) Now what about when you multiply something like this: \[(2\sqrt{3})(6\sqrt{2})\]

OpenStudy (cggurumanjunath):

12*SQRT(6)

OpenStudy (cggurumanjunath):